OpenStudy (anonymous):

Find the equation of the plane in R3 that contains the point A(-6, -2, -2), and is perpendicular to the line through to points (0, -2, -6) and (2, -4, -2

OpenStudy (anonymous):

To find the equation of plane , you need a point and normal vector point A(-6, -2, -2) normal vector from (0, -2, -6) to (2, -4, -2) <2,-2,4>

OpenStudy (amistre64):

the line given is the normal right?

OpenStudy (anonymous):

a(x+6)+b(y+2)+c(z+2)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

<2,-2,4> could be scaled down to <1,-1,2>

OpenStudy (amistre64):

(0, -2, -6) and (2, -4, -2) +2 +6 +2 +6 ------------------------ ( 0 , 0 , 0) (2,-2,4) .... yep good eye

OpenStudy (anonymous):

how you calculate the normal..?

OpenStudy (anonymous):

I believe we just did 1(x+6)-1(y+2)+2(z+2)

OpenStudy (amistre64):

the normal is perpendicular to the plane; so you just need to know the vector that points from the 2 other given points since that IS perpendicular to the plane

OpenStudy (anonymous):

and it's equal to zero..?

OpenStudy (amistre64):

a little more specific please

OpenStudy (anonymous):

confused.. may I noe the formula for the normal?

OpenStudy (anonymous):

1(x+6)-1(y+2)+2(z+2)=0 x-y+2 z=-8

OpenStudy (amistre64):

subtract one point from the other that constitutes the perpendicular line to the plane as stated in the problem

OpenStudy (amistre64):

point1 = (0, -2, -6) -point2 = -(2, -4, -2) --------------------- normal = <-2, 2,-4>, or its negation

OpenStudy (amistre64):

this line is parallel to the normal of your plane

OpenStudy (anonymous):

how you decide if the line is parallel or perpendicular?

OpenStudy (amistre64):

..... i used the information given inthe stated problem.

OpenStudy (anonymous):

ok. let me think about it for awhile.. :)

OpenStudy (amistre64):

"...and is perpendicular to the line through to points (0, -2, -6) and (2, -4, -2)" means: whose normal vector is parallel to the line through to points (0, -2, -6) and (2, -4, -2)

OpenStudy (anonymous):

:) ah. okay! get it now.. :))

OpenStudy (amistre64):

good :)

OpenStudy (amistre64):

once you know the normal; its just the vector that directs you from one point to the next, you can apply it to the equation of a plane with the given point

OpenStudy (amistre64):

right?

OpenStudy (anonymous):

yup.

OpenStudy (anonymous):

i get the answer now..

OpenStudy (anonymous):

Find the equation of the plane in R3 that contains the point A(2, 5, 5), and is perpendicular to the line x = 6 - 3 t y = -4 - 3 t z = -3 t

OpenStudy (anonymous):

do u noe how to do this..? im stuck for looking the value for the normal..

OpenStudy (amistre64):

the line they give you is in parametric form; which means that it has a point and a vector hidden within the information itself. And its the vector thats hidden in there that we want to use.

OpenStudy (amistre64):

remember that a line can be form by starting at a given point and scaling a vector to any length. In other words, the constant parts are the point components, and the coefficients of the variables are actually are vector components. the point given in the line is: (6,-4,-3) the vector attached to that point is: <-3t,-3t,-3t>, or simply -3t<1,1,1> our normal to the plane for the question is then <1,1,1>